# A Calculus of Overlays

### v1

@@ 6. Discussion
6.1. Relationship to $\lambda$-Calculus
Three facts establish the relationship between Overlay-Calculus and the $\lambda$-calculus.
More elementary and stronger than Scott domains.
The semantics induced by the translation (Section 4) requires only powerset lattices and the Knaster–Tarski fixed-point theorem—no function spaces, no inverse limits, no Scott continuity. Yet this semantics is fully abstract (Theorem 4.3), whereas Scott’s $D_{\infty}$ model (Scott, 1976; Wadsworth, 1976) is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
$\lambda$-Calculus embeds into Overlay-Calculus.
The five-rule translation of Section 4 maps every ANF $\lambda$-term to an overlay tree. The translation uses only single-target qualified this—the full multi-target semantics is not needed to embed $\lambda$-calculus.
Multi-target qualified this is the natural primitive.
If the semantic functions of Section 3 were restricted to single-target resolution (as in DOT’s single-valued self variable (Amin et al., 2016)), the formulas would need to reject various structurally legitimate cases—adding complexity, not removing it. Multi-target qualified this is not an artificial extension; it is what falls out when no restrictions are imposed. The $\lambda$-calculus translation uses only single-target qualified this, suggesting that $\lambda$-calculus itself represents an artificially restricted fragment of the natural model. Appendix H presents a detailed analysis of the asymmetry in problem difficulty between the two systems.
6.2. Observed Phenomena
Several properties emerge in practice without having been designed:
  • The Expression Problem does not arise (Section 5, Appendix H).
  • Programs are CPS-agnostic: empty record slots serve as dependency injection points, making the same program text CPS or direct-style depending on assembly (Appendix G).
  • Records are tries, and inheritance is trie union—an immutable analogue of random-access memory (Appendix A).
  • Unlike SKI combinators (Schönfinkel, 1924; Curry and Feys, 1958), which are isomorphic to $\lambda$-calculus in problem difficulty, Overlay-Calculus exhibits genuine asymmetry (Section 4.1).
6.3. Properties of Inheritance
Because records are sets, inheritance is inherently commutative and idempotent, making Overlay-Calculus transparent to linearization order. Deduplication of inherited overlays is purely an optimization.
6.4. Relationship to Practical Systems
Several classes of declarative systems that have emerged independently in practice turn out to be instances of Overlay-Calculus.
Configuration and object-oriented languages
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024)—whose recursive attribute merging motivated this work—can be understood as an implementation of Overlay-Calculus with additional features (type checking, error reporting, FFI to the Nix language). Detailed comparisons with CUE, Jsonnet, Dhall, with mixin and trait calculi, and with the NixOS module system’s lazy evaluation limitations under multiple inheritance routes are given in Section 7.
Effect systems
Algebraic effects with handlers (Plotkin and Power, 2003; Pretnar, 2015) and monad transformers structure computational effects through composition. The CPS-agnostic property (Appendix G) shows how Overlay-Calculus naturally supports effect handler composition: empty API slots can be filled with effect handlers (CPS interpretation) or direct implementations (direct-style interpretation). The same pattern appears in algebraic effect systems, without explicit monad machinery.
@@ 6.4. Relationship to Practical Systems
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles Overlay-Calculus—later layers override earlier layers, files from all layers remain accessible—but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented Overlay-Calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance—a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
7. Related Work
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction. Cook’s key insight—that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods—is one of the starting points of the present work. The model has four primitives (records, functions, generators, wrappers) and requires a fixed-point combinator; composition is asymmetric (a wrapper modifies a generator, not vice versa). Whether the same insight can be captured with fewer primitives and symmetric inheritance was a natural question.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass—unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity was worth examining. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure—composing two mixins that define the same method name results in an override, not a union mount. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them—without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a significant improvement over mixins. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than opportunities for deeper merging. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope—a common situation under inheritance—single-valued resolution must choose one route and discard the others, losing information. Whether qualified this can resolve to a set of paths, letting all routes contribute equally, had not been explored.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects—rather than classes—as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions (taking self as a parameter), and object extension is asymmetric (method override replaces the previous definition). Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language—for instance, through coalgebraic observation rather than reduction—was a question these calculi left open.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined (a type error); the calculus avoids same-label composition rather than giving it meaning. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records (no nesting) and treat records as data (initial algebras): a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation—the properties needed for records to serve as a computational substrate—are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix C). DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics (unifying incompatible scalars yields $\bot$); whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative (the right-hand side wins on scalar conflicts), and deep merging requires explicit opt-in (+: syntax) on a per-field basis—so union mount semantics is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and union mount semantics
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override—the same union mount semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system got nearly all of the high-level design right: union mount inheritance, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. Its practical success is remarkable—the same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option—the multi-path situation that recursive merging naturally gives rise to—which motivated the coalgebraic semantics of Section 3.
SKI combinator calculus
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus can be reformulated without bound variables using three combinators (S, K, I). SKI is function-free at the object level and Turing complete. However, bracket abstraction translates between SKI and the $\lambda$-calculus preserving computational structure, so the two systems are isomorphic in terms of problem difficulty (Section 4): what is hard in one remains hard in the other. Whether a function-free calculus can exhibit a genuine shift in problem difficulty—making some problems structurally easier than in the $\lambda$-calculus, rather than merely encoding them differently—is the question that motivates the comparison in Section 4.

### v2

@@ 6. Discussion
6.1. Emergent Phenomena
Several features of overlay-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
  1. (1)
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
  3. (3)
    Commutative, idempotent merge. Inheritance in overlay-calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
  4. (4)
    Relational semantics. Every overlay-calculus operation is defined by recursion on overlay trees, not on individual values. When an overlay tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued overlays $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
  5. (5)
    Function color blindness. We implemented a web application as overlay files (Yang, 2025b), initially targeting Python’s synchronous sqlite3 and http.server modules. The business logic overlay file contained no async-specific code. Later, we replaced the FFI layer with one built on aiosqlite and starlette; the business logic overlay file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the overlay file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
Each of these features is primitive in overlay-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms.
6.2. Classification of Semantic Variants
The goals of declarative programming rule out several properties of the model: separate primitives for methods, classes, types, and objects; mutable reference cells; domain-theoretic approximation; and asymmetric override. Ruling these out leaves a minimal syntax of three constructs: records, definitions, and inheritance.
Two degrees of freedom remain. These dimensions are independent, and each combination yields a coherent semantics. They arise in any system whose definitions form a tree of named scopes, where each node is a path and references are path-dependent, even in systems ruled out above, such as Scala, which has separate primitives for methods, classes, objects, and types.
Existing designs occupy several of the resulting cells:
No deep merge Deep merge
Early binding CUE (van Lohuizen, 2019)
Dynamic scope union file systems, Jsonnet (Cunningham, 2014)
Lexical scope Scala (Odersky et al., 2004) NixOS modules (NixOS Contributors, 2024)
Multi-target $\mathtt{this}$ overlay-calculus
Early binding
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a field in a nested record cannot refer back to properties of its enclosing context contributed by later composition. Early binding combined with deep merge is coherent and practically useful, but cross-overlay cooperation requires explicit wiring rather than inheritance: a definition cannot observe properties contributed by overlays composed after it.
Dynamic scope
With dynamic scope, references resolve against the context at the point of use rather than the definition site. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. When the Scala multi-path example (Appendix B) is translated to Jsonnet, it does not error, but outer resolves to the merged MyObjectA rather than to Object1 or Object2 separately, silently discarding the distinction between the two inheritance routes. Dynamic scope is flexible but loses the ability to track which inheritance route a reference came from.
Deep merge
Without the second case of equation (3), $\mathtt{overlays}(p)$ would always return $\{p\}$, a single element. When two independently authored overlays both define a label with the same name at the same scope level, only the definition at the literal path $p$ would contribute to $\mathtt{supers}$ and hence to $\mathtt{properties}$; the other definitions would be silently invisible. This is the situation in most composition systems: pytest permits at most one fixture per name per scope; trait systems (Schärli et al., 2003) flag same-name methods as conflicts. Under any of these rules, two independently authored overlays that both extend an entity with the same name, which is the situation that arises naturally when solving the Expression Problem (Wadler, 1998), cannot be composed. The non-trivial case of equation (3) removes this limitation: it collects all paths that share the same label and merges their definitions rather than shadowing each other. This is the record-level counterpart of method override in object-oriented languages, and the mechanism by which the calculus admits extensibility in both dimensions of the Expression Problem (Wadler, 1998) (Section 5).
Multi-target qualified this
The qualified this reference $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Scheme (Sussman and Steele Jr., 1975) introduced lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix B) and the NixOS module system rejects the duplicate option declaration (Appendix B); neither can express that all routes are equally valid.
Overlay-calculus combines the predictability of lexical scope with the flexibility of dynamic scope. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural, not incidental: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope. If the equations were restricted to single-target resolution, as in DOT’s single-valued self variable (Amin et al., 2016), they would need to reject structurally legitimate cases, adding complexity rather than removing it.
Consequences of the alternative choices
Choosing differently in any dimension requires introducing machinery that is absent from the equations of Section 3. With early binding, the equations must freeze reference targets at definition time, so a definition cannot observe properties contributed by later-composed overlays. With dynamic scope, the equations must resolve references against a runtime context rather than the definition-site structure, losing the ability to distinguish inheritance routes. Restricting $\mathtt{overlays}(p)$ to $\{p\}$ means same-label definitions from different inheritance sources no longer share identity, requiring either a method resolution order (MRO) (Barrett et al., 1996) to linearize them or a shadowing rule to select one and discard the rest, which is why systems without deep merge rarely support path-dependent multiple inheritance in practice. Restricting $\mathtt{this}$ to a single path requires a mechanism to select among inheritance routes (Scala rejects the ambiguity at compile time; the NixOS module system raises a duplicate-declaration error), and loses contributions from all non-selected routes. These mechanisms, including early-binding freezing, runtime context threading, MRO, shadowing, route selection, and error reporting, are more artificial than their absence. The equations of Section 3 contain none of them.
7. Related Work
Minimal Turing-complete models
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators, which are functions from self to complete object, and wrappers, which are functions that modify generators, solved by a fixed-point construction. Cook’s key insight, that inheritance is a general mechanism applicable to any form of recursive definition and not only object-oriented methods, is one of the starting points of the present work. The model has four primitives—records, functions, generators, and wrappers—and requires a fixed-point combinator; composition is asymmetric, as a wrapper modifies a generator but not vice versa.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass, unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity had not been established. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure: composing two mixins that define the same method name results in an override, not deep merge. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them; without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a departure from mixin ordering. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than recursive merges. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope, a common situation under inheritance, single-valued resolution must choose one route and discard the others, losing information. The framework provides no mechanism for set-valued resolution.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects rather than classes as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions that take self as a parameter, and object extension is asymmetric since method override replaces the previous definition. Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language, for instance through coalgebraic observation rather than reduction, was not explored in these calculi.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined and raises a type error; the calculus leaves same-label composition undefined. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records with no nesting and treat records as data, that is, initial algebras: a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation, which are the properties needed for records to serve as a computational substrate, are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B). DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules. Whether these three roles could be unified under a single construct remains an open question in the DOT line of work. DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and deep merge
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same overlay semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the coalgebraic semantics of Section 3.
@@ 7. Related Work
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles overlay-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented overlay-calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.